A body travels 200 cm in the first 2 sec and 220cm in next 5 sec. Calculate the velocity at the end of the seventh second from the start
 

Grade:10

1 Answers

Arun
5 years ago

Let u, v & w be the initial, at the end of 2s & at the end of 6 s respectively.

Using average velocity *time = distance

(u + v )*2/2 = 200 we get u+ v = 200 m/s

Similarly, v + w =2 *220/4 = 110 m/s

Acceleration = ( w-u)/ time = ( 110 - 200)/ (4+2) = -15 m/s^2

We have u+ v = 200 m/s => 2u + at = 2u - 30 => u = 115 m/s.

Velocity at the end of 7 s is u + 7a = 115 - 105 = 10 m/s.